OpenStudy (anonymous):

If m ≠ 1, then m/m-1>1

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{ same }}{\cancel{ same }}=1\qquad thus \qquad\cfrac{\cancel{ same }}{\cancel{ same }}-1=?\)

OpenStudy (anonymous):

I think @cpeters3535 meant \[\frac{m}{m-1}>1\] which is not true for all \(m<1\). Consider \(m=\dfrac{1}{2}\), then \[\frac{\frac{1}{2}}{\frac{1}{2}-1}=\frac{\frac{1}{2}}{-\frac{1}{2}}=-1\not{>}1\]